# Estimating angular momenta of fission fragments from isomeric yield ratios using TALYS

### 22 Feb 2024 v1

@@ II.3 Angular momentum
As seen in Figure 5, the calculated IYRs from different initial FFs show a similar dependency on the $J_{rms}$, indicating that common factors affect the calculated IYR. One such factor is the level scheme, including the two long-lived states, of ${}^{131}$e. On the other hand, the observed variations between the curves could indicate that the FFs have different impacts on IYR.
Considering all contributing FFs, one can calculate the weighted average IYR as a function of $J_{rms}$
@@
As an example, the calculated average IYRs of the FP ${}^{131}$e are presented in black points in Figure 5. As seen, the average IYR matches the experimental ratio at a $J_{rms}$ of $12.3~{}\hbar$. Furthermore, the calculated IYRs cross the experimental ratio minus (plus) one sigma at a $J_{rms}$ of $11.5~{}\hbar$ ($13.4~{}\hbar$). Hence, a value of $12.3^{1.1}_{0.8}~{}\hbar$ is assigned as the most likely $J_{rms}$ of the FFs de-exciting to the FP ${}^{131}$e.
@@
The procedure described above is repeated for all 31 IYRs. However, in some cases the calculations do not reproduce the experimental value for any $J_{rms}$ in the range 0.2 to 25 $\hbar$. Figure 6 shows the example of the IYR of the FP ${}^{121}$d, obtained from the de-excitations of the FFs ${}^{121-129}$d*. As seen, the average IYRs do not reach the experimental ratio for any $J_{rms}$, although some of the ratios for single FFs ($A={126-129}$) reproduce the experimental ratio.
@@
Figure 6: Calculated IYRs of the FP ${}^{121}$d from all excited FFs, ${}^{121-129}$d*, contributing to the production of ${}^{121}$d. The horizontal line and grey region represent the measured IYR with uncertainties of the FP ${}^{121}$d [21].
@@
If we assume that the experimentally determined IYR is correct, there are at least three other possible reasons for the failure of the model to reproduce the experimental data. First of all, besides $J_{rms}$, the values of the other parameters ($C$, $\bar{E}_{offset}$, $\bar{\sigma_{E}}$) also affect the energy and angular momentum population of the FFs. In principle, also these parameters could be optimized against experimental data. However, this would require a much larger data set, probably including neutron and gamma multiplicities, in order not to have an underdetermined problem. Secondly, the de-excitation process, as described by the Hauser-Feshbach approach in TALYS, might need adjustment. Thirdly, and probably the most likely reason, is the lack of experimental excitation levels, and poor knowledge of de-excitation paths toward the long-lived states of the FP. In the case of ${}^{121}$d, the number of known levels in the complete scheme is 9, indicating a poor knowledge of the nuclear structure.
@@
Furthermore, several studies [29, 30, 31, 32] have shown that a parameter (rspincut) in TALYS, as a multiplication to the spin cut-off parameter, can have a substantial impact on the calculated isomeric yield ratios for light particle-induced reactions. This parameter is used to adjust the predicted spin distribution from the level density models before applying it to the angular momentum distribution. In these studies, the suggested value of the rspincut parameter is about 0.5, which reduces the width of angular momentum distribution. This value was tested with the surrogate model, but no significant change could be observed. This can be understood, as in default TALYS calculations, the level density model and rspincut determine the angular momentum distribution of the excited nucleus formed in the reactions, while in our surrogate model, it is determined by the $Ex-J$ matrix, and hence independent of the parameters in TALYS.
II.4 Primary contributor
@@
As seen in Figure 5 and 6, the IYR as a function of $J_{rms}$ behaves differently for different fission fragments. In forming the weighted average ratios in Equation 5 the relative contribution from each fragment is taken into account. Although all possible contributors are included, the ratio is largely determined by two to four of the contributors. This can be shown by studying the relative contributions, which can be calculated by multiplying the independent yield of each fragment with the fraction of that fragment that decays to the studied FP. The result for ${}^{131}$e and ${}^{121}$d is presented in Table 2. At the same time, the average number of neutrons emitted from these fragments de-exciting to a specific fission product can be calculated.
@@
Table 2: Contributions to the production of ${}^{131}$e and ${}^{121}$d in the TALYS calculation. Contributions larger then 10 % are marked in bold font.
@@ III Results and discussion
When experimental determinations of angular momentum in fission are discussed in the literature, the deduced values are usually associated with the fission products in question, even though it is the angular momentum of the initial fragment that is sought. In this study, we are also unable to associate an estimated $J_{rms}$ to a specific FF, but rather to a range of FF isotopes. With Table 2 in mind, the deduced value of $12.3^{1.1}_{0.8}~{}\hbar$ can for example be considered a reasonable estimate of the $J_{rms}$ for FFs of Tellurium isotopes in the mass range $132-135$.
@@
As mentioned in Section II.2 De-excitation with TALYS, there is a limit in TALYS of 30 discrete spin values. Hence, if a FF is initialized with a $J_{rms}$ larger than about $20~{}\hbar$ the angular momentum distribution will be cut. As a result, values of $J_{rms}$ larger than $20~{}\hbar$ should be considered less trustworthy.
@@ III.2 Average angular momenta
J. Wilson et al. have suggested that the angular momenta of fission fragments are generated from the collective motion of nucleons in the ruptured neck of the fissioning system [15]. Based on the Rayleigh distribution and the Fermi gas model, an equation to describe the observed mass dependency of the average angular momentum is proposed:
@@
in which $c$ is a free parameter and $A_{F}$ is the fragment mass. $A_{N}$ is a parameter that for fission fragments near the doubly magic ${}^{132}$n is $A_{F}-130$, while for light fragments near the doubly magic ${}^{78}$i it is $A_{F}-78$. In their paper, J. Wilson et al. fit this equation to their data for fast (average energy 1.9 MeV) neutron-induced fission of ${}^{238}$. However, they do not present any data, and there is also no description or prediction of the $J^{av}(A)$ in the symmetry region.
@@
Equation (6) was also fitted to the proton-induced data for A larger than 131, and is shown as a red curve in Figure 8. The model describes the data rather well. However, the average angular momenta associated to ${}^{130}$n and ${}^{131}$e cannot be described by Equation (6). The failure to describe these nuclides might indicate that Eq. 6 is not valid for mass numbers close to 130. This observation is also supported by the fact that the model predicts $J^{av}$ to be zero for A = 130. On the other hand, it could be argued that the minimum angular momentum should be observed for the most spherical fission products around the doubly magic ${}^{132}$n.
Figure 8: Average angular momentum as a function of mass number of the resulting fission product. The colored points present data from proton-induced fission of ${}^{238}$, and the red curve is fit to these data for A greater than 131. The red dashed line is obtained from a linear fit to the data for A less than 133. The black points and curves present the average angular momenta for the fission fragments in the fast neutron-induced fission of ${}^{238}$, and a fit to the data from J. Wilson et al. [15].
J. Wilson et al. do not present any data for the angular momentum of fission products in the symmetric mass region (105 $\leq$ A $<$ 130). Here our data reveal that the average angular momenta decrease with increasing mass number in contrast with the predicted mass dependency by J. Wilson. In Wilson’s model, the average momentum is generated from the collective motion of nucleons in the ruptured neck, which always increases from the starting mass number (78 or 130 or any other values). Hence, it is impossible to predict a decreasing trend of the average angular momentum with Wilson’s model. There is a need for a different mechanism to explain the observed decreasing trend of the average angular momentum in the symmetry region.
It might be worth mentioning that a linear fit to the data gives an average decrease by $0.3\pm 0.2~{}\hbar$ per atomic mass unit, though there is not a reasonable guess on what it means by now.
@@ III.3 Neutron emission
From the TALYS output of the surrogate model, information on the neutron multiplicity for each fission fragment can be extracted. From this information, the average neutron multiplicity for each fragment decaying to a certain fission product can be derived. In figure 9 this is plotted together with the same information obtained directly from the GEF model.
Figure 9: The average number of neutrons emitted from the primary fission fragments de-exciting to a specific fission product as derived from TALYS, compared with the same information from GEF.
In Figure 9, a similar behaviour of the two models is observed. Although the neutron multiplicity in proton-induced fission in GEF is not validated using experimental data and should hence not be taken as a benchmark, the agreement builds some confidence in the surrogate model. The discrepancies that can be observed are likely due to the different ways of dealing with the de-excitation process and different assumptions of angular momenta of the fission fragments.
@@ III.4 Angular momentum of the primary fission fragment
Considering that the derived angular momentum is a property of the initial fragments, it would be interesting to study how this varies with, for example, the fragment mass. Using the neutron emission information of the de-excitation from TALYS, shown in Figure 9, the average mass number of the fission fragments de-exciting to a certain fission product was calculated. As a result, the fission product mass number in Figure 8 is shifted to the most probable FF mass number, as presented in Figure 10.
Figure 10: The average angular momentum as a function of the average mass number of the primary fission fragments de-exciting to the observed fission products. The dotted line represents a fitting to the angular momentum as a function of the primary fission fragments in the symmetry region.
Assuming the same mass dependence of the angular momentum of the fission fragments as for the fission products, the data in Figure 10 has been fitted with Equation 6 for A $>$ 134 and with a straight line in the symmetric mass region. The decreasing trend of the average angular momentum for A $<$ 134 is similar to that for A $<$ 132 in Figure 8. However, the minimum of the average angular momenta is shifted by approximately two mass units towards higher masses. This shift makes sense, considering that the neutron multiplicity in the proton-induced fission of ${}^{238}$ at 20 MeV is about 2 [33], which is similar to the fissioning system in this study.
@@ IV Conclusion and outlook
A surrogate model of GEF that can be used to generate energy vs angular momentum matrices describing excited fission fragments has been developed. Using TALYS to calculate the de-excitation of the fragments while varying the parameters of the model, the impact of the state of primary FFs on the IYRs in nuclear fission has been investigated.
31 IYRs in proton-induced fission of ${}^{238}$ at 25 MeV are compared with the calculated ratios from the model. Based on that comparison, 24 estimates of the root-mean-square angular momentum, $J_{rms}$, have been estimated. For some fission products, the estimated $J_{rms}$ are compared with the values derived by V. Rakopoulos, and differences between the two methods are discussed.
It has been shown that the mass dependence of average angular momentum $J^{av}$ for the nuclides with mass numbers larger than 131 could be reasonably well described by the model suggested by J. Wilson et al. [15]. This dependence is based on the assumption that the angular momentum is generated from the collective motion of nucleons in the ruptured neck of the fissioning system.
In the symmetric mass region, a decreasing trend of the average angular momentum of the fission products is observed for the first time. This observation could not be explained by J. Wilson’s proposal, and needs a different mechanism to model the angular momentum generation in the symmetry region.
To link the average angular momentum to the primary fission fragments, the number of emitted neutrons is calculated for each fragment. This way, the average angular momentum as a function of the average mass number of the primary fission fragment is obtained. In this description, the minimum of the angular momentum is close to mass number 134, deviating from the shell closure at 132 by two units.
Considering the fact that the surrogate model fails to reproduce the experimental data in some cases shows that the model needs further investigation and development. In particular, the impact of the (un)known levels scheme of the fission products, and possible dependencies of the IYR on parameters other than the angular momentum, are worth future studies. Also, the effect of other parameters in TALYS, such as rspincut possibly related to the momentum of inertia [26], should be investigated in order to improve the IYR calculation.
@@ V Acknowledgement
This work was supported by the Swedish research council Vetenskapsrådet (Ref. No. 2017-06481), the European Commission within the Seventh Framework Programme through Fission-2013-CHANDA (Project No. 605203), the Swedish Radiation Safety Authority (SSM).

### 18 Mar 2024 v3

@@ II.3 Angular momentum
As seen in Figure 5, the calculated IYRs from different initial FFs show a similar dependency on the $J_{rms}$, indicating that common factors affect the calculated IYR. One such factor is the level scheme, including the two long-lived states, of ${}^{131}$e. On the other hand, the observed variations between the curves could indicate that the FFs have different impacts on the IYR.
Considering all contributing FFs, the weighted average IYR as a function of $J_{rms}$ was calculated,
@@
As an example, the calculated average IYRs of the FP ${}^{131}$e are presented in black points in Figure 5. As seen, the average IYR matches the experimental ratio at a $J_{rms}$ of $12.3~{}\hbar$. Furthermore, the calculated IYRs cross the experimental ratio minus (plus) one sigma at a $J_{rms}$ of $11.5~{}\hbar$ ($13.4~{}\hbar$). Hence, an average value of $12.3^{1.1}_{0.8}~{}\hbar$ is assigned as the most likely $J_{rms}$ of the FFs de-exciting to the FP ${}^{131}$e.
@@
The procedure described above is repeated for all 31 IYRs. However, in some cases the calculations do not reproduce the experimental value for any $J_{rms}$ in the range 0.2 to 25 $\hbar$. Figure 6 shows the example of the IYR of the FP ${}^{121}$d, obtained from the de-excitations of the FFs ${}^{121-129}$d*. As seen, the average IYRs do not reach the experimental ratio for any $J_{rms}$, although some of the ratios for single FFs ($A={126-129}$) can reproduce the experimental ratio.
@@
Figure 6: Calculated IYRs of the FP ${}^{121}$d from all excited FFs, ${}^{121-129}$d*, contributing to the production of ${}^{121}$d. The horizontal line and grey region represent the measured IYR of the FP ${}^{121}$d with uncertainties [24].
@@
If we assume that the experimentally determined IYR is correct, there are at least three other possible reasons for the failure of the model to reproduce the experimental data. First of all, besides $J_{rms}$, the values of the other parameters ($C$, $\bar{E}_{offset}$, $\bar{\sigma_{E}}$) also affect the energy and angular momentum population of the FFs. In principle, also these parameters could be optimized against experimental data. However, this would require a much larger data set, probably including neutron and gamma multiplicities, in order not to have an underdetermined problem. Secondly, the de-excitation process, as described by the Hauser-Feshbach approach in TALYS, might need adjustment. Thirdly, and probably the most likely reason, is the lack of experimental excitation levels, and poor knowledge of de-excitation paths toward the long-lived states of the FP. In the case of ${}^{121}$d, the number of known levels in the complete scheme is only 9, indicating a poor knowledge of the nuclear structure.
@@
Furthermore, several studies [32, 33, 34, 35] have shown that the width of the angular momentum distribution can have a substantial impact on the calculated isomeric yield ratios for light particle-induced reactions. In TALYS, the width of the predicted spin distribution from the level density models can be controlled by scaling the spin cut-off parameter using a parameter called Rspincut. The studies indicate that the optimal value of the Rspincut parameter in light particle-induced reactions is about 0.5, which reduces the width of angular momentum distribution compared to the default value of 1. In this study, this value was tested with the surrogate model, leading to, in most cases, a significantly larger angular momentum than that obtained with an Rspincut of 1. However, changing Rspincut to 0.5 does not improve the surrogate model’s ability to reproduce the experimentally determined IYRs. Since many parameters in the model affect the calculated IYR, and most of them are not yet optimized due to lack of data, the impact of Rspincut for this data set could not be determined. Hence, only the results with the default value of Rspincut of 1 are presented in this paper.
II.4 Primary contributors
@@
As seen in Figure 5 and 6, the IYR as a function of $J_{rms}$ behaves differently for different FFs. In forming the weighted average ratios in Eq. (5) the relative contribution from each fragment is taken into account. Although all possible contributors are included, the ratio is largely determined by two to four of the contributors. This can be shown by studying the relative contributions, which can be calculated by multiplying the independent yield of each fragment with the fraction of that fragment that decays to the studied FP. The result for ${}^{131}$e and ${}^{121}$d is presented in Table 2. At the same time, the average number of neutrons emitted from these fragments, when de-exciting to a specific FP, can be calculated.
@@
Table 2: Contributions to the production of ${}^{131}$e and ${}^{121}$d in the TALYS calculations. Contributions larger then 10 % are marked in bold font.
@@ III Results and discussion
When experimental determinations of angular momentum in fission are discussed in the literature, the deduced values are usually associated with the FPs in question, even though it is the angular momentum of the initial fragment that is sought. In this study, we are also unable to associate an estimated $J_{rms}$ to a specific FF, but rather to a range of FF isotopes. With Table 2 in mind, the deduced value of $12.3^{1.1}_{0.8}~{}\hbar$ can for example be considered a reasonable estimate of the $J_{rms}$ for FFs of Tellurium isotopes in the mass range $132-135$.
@@
As mentioned in Section II.2, there is a limit in TALYS of 30 discrete spin values. Hence, if a FF is initialized with a $J_{rms}$ larger than about $20~{}\hbar$ the angular momentum distribution will be significantly cut. As a result, values of $J_{rms}$ larger than $20~{}\hbar$ should be considered less trustworthy.
@@ III.2 Average angular momenta
J. Wilson et al. have suggested that the angular momenta of FFs are generated from the collective motion of nucleons in the ruptured neck of the fissioning system [15]. Based on the Rayleigh distribution and the Fermi gas model, an equation to describe the observed mass dependency of the average angular momentum is proposed:
@@
in which $c$ is a free parameter and $A_{F}$ is the fragment mass. $A_{N}$ is a parameter that for FFs near the doubly magic ${}^{132}$n is $A_{F}-130$, while for light fragments near the doubly magic ${}^{78}$i it is $A_{F}-78$. In their paper, J. Wilson et al. fit this equation to their data for fast (average energy 1.9 MeV) neutron-induced fission of ${}^{238}$. However, they do not present any data, and there is also no description or prediction of the $J^{av}(A)$, in the symmetric mass region.
@@
Eq. (6) was also fitted to the proton-induced data for A larger than 131, and is shown as a red curve in Figure 8. The model describes the data rather well. However, the average angular momenta associated to ${}^{130}$n and ${}^{131}$e are much higher than the values predicted by Eq. (6). The failure to describe these nuclides might indicate that Eq. (6) is not valid for mass numbers close to 130. This observation is also supported by the fact that the model predicts $J^{av}$ to be zero for A = 130. On the other hand, it could be argued that the minimum angular momentum should be observed for the most spherical FPs around the doubly magic ${}^{132}$n.
Figure 8: Average angular momentum as a function of mass number of the resulting FP. The colored points present data from proton-induced fission of ${}^{238}$, and the red curve is fit to these data for A greater than 131. The red dashed line is obtained from a linear fit to the data for A less than 133. The black points and curves present the average angular momenta for the FFs in the fast neutron-induced fission of ${}^{238}$, and a fit to the data from J. Wilson et al. [15].
J. Wilson et al. do not present any data for the angular momentum of FPs in the symmetric mass region in which mass numbers are close to half of the fissioning nucleus. Our data, however, suggest that the average angular momenta in this region decrease with increasing mass number. A linear fit to the data gives an average decrease by $0.3\pm 0.2~{}\hbar$ per atomic mass unit. This behaviour can not be explained with the model presented by J. Wilson et al. Since the average angular momentum is assumed to be generated from the collective motion of nucleons in the ruptured neck, the angular momentum must increase with the number of nucleons. Even if one changes the parameters in Eq. (6), it is impossible to predict a decreasing trend of the average angular momentum. Hence, a different mechanism is needed to explain the observed decreasing trend of the average angular momentum in the symmetry region.
@@ III.3 Neutron emission
From the TALYS output of the surrogate model, information on the neutron multiplicity for each FF can be extracted. From this information, the average neutron multiplicity for each fragment decaying to a certain FP can be derived. In Figure 9, this is plotted together with the same information obtained directly from the GEF model, and a similar behaviour of the two models is observed. Although the neutron multiplicity in proton-induced fission in GEF is not validated against experimental data, and should hence not be taken as a benchmark, the agreement builds some confidence in the surrogate model. The discrepancies that can be observed are likely due to the different ways of dealing with the de-excitation process, and different assumptions of angular momenta of the FFs.
Figure 9: The average number of neutrons emitted from the primary FFs de-exciting to a specific FP as derived from TALYS, compared with the same information from GEF.
@@ III.4 Angular momentum of the primary fission fragment
Considering that the derived angular momentum is a property of the initial fragments, rather than the FPs, it would be interesting to study how this varies with the fragment mass. Using the neutron emission information of the de-excitation from TALYS, shown in Figure 9, the average mass number of the FFs de-exciting to a certain FP was calculated. In this way, the fission product mass number in Figure 8 can be shifted to the most probable FF mass number, and the result is presented in Figure 10.
Figure 10: The average angular momentum as a function of the average mass number of the primary FFs de-exciting to the observed FPs. The dashed line represents a linear fit to the angular momentum as a function of the primary FFs in the symmetry region. The red curve is a fit of Eq. (6) to the angular momenta as a function of the primary FFs for A$>$ 134.
Based on the assumption by Wilson et al., it makes more sense to fit the resulting parametrization (Eq. (6)) as a function of the mass of the primary fragments. The shift of the minimum of the average angular momenta from 132 to 134 in Figure 10, implies that the FFs around mass number 134 have the lowest angular momenta. Considering that the neutron multiplicity for FFs around A= 134 in proton-induced fission of ${}^{238}$ at 20 MeV is about 2 [36], these FFs will de-excite to FPs around mass number 132.
The decreasing trend of the average angular momentum for A $<$ 134 in Figure 10 is similar to that for A $<$ 132 in Figure 8. This observation supports our discussion in section III.2.
@@ IV Conclusion and outlook
A surrogate model of GEF that can be used to generate energy vs angular momentum matrices describing excited FFs has been developed. Using TALYS to calculate the de-excitation of the fragments while varying the parameters of the model, the impact of the state of primary FFs on the IYRs in nuclear fission has been investigated.
31 IYRs in proton-induced fission of ${}^{238}$ at 25 MeV are compared with the calculated ratios from the model. Based on that comparison, the root-mean-square angular momentum, $J_{rms}$, of 24 FPs have been estimated. When applicable, the estimated $J_{rms}$ are compared with the values derived by V. Rakopoulos et al., and differences between the two methods are discussed.
It has been shown that the mass dependence of the average angular momentum $J^{av}$ for the nuclides with mass numbers larger than 131 could be reasonably well described by the model suggested by J. Wilson et al. [15]. This dependence is based on the assumption that the angular momentum is generated from the collective motion of nucleons in the ruptured neck of the fissioning system.
In the symmetric mass region, a decreasing trend of the average angular momentum of the FPs is observed for the first time. This observation could not be explained by the model proposed in Wilson’s article and needs a different mechanism to model the angular momentum generation in the symmetry region.
To link the average angular momentum to the primary FFs, the number of emitted neutrons is calculated for each fragment. This way, the average angular momentum as a function of the average mass number of the primary FF is obtained. In this description, the minimum of the angular momentum is close to mass number 134, deviating from the shell closure at 132 by two mass units.
The fact that the surrogate model fails to reproduce the experimental data in some cases shows that the model needs further investigation and development. In particular, the impacts of the (un)known level scheme of the FPs, and possible dependencies of the IYR on parameters other than the angular momentum, are worth future studies. Also, the effect of other parameters in TALYS, such as Rspincut possibly related to the momentum of inertia [29], should be investigated in order to improve the IYR calculation.
@@ V Acknowledgement
This work was supported by the Swedish research council Vetenskapsrådet (Ref. No. 2017-06481), the European Commission within the Seventh Framework Programme through Fission-2013-CHANDA (Project No. 605203), the Swedish Radiation Safety Authority (SSM). The authors would like to acknowledge Harald och Louise Ekmans forskningsstiftelse for the generous grant at Sigtunastiftelsen.